### v2

# Leray-Schauder Mappings for Operator Learning

Abstract.
We present an algorithm for learning operators between Banach spaces, based on the use of Leray-Schauder mappings to learn a finite-dimensional approximation of compact subspaces. We show that the resulting method is a universal approximator of (possibly nonlinear) operators. We demonstrate the efficiency of the approach on two benchmark datasets showing it achieves results comparable to state of the art models.
Keywords: Universal approximation; nonlinear operator; Burgers’ equation; integral equations.
1. Introduction
Operator learning is a branch of deep learning involved with approximating (potentially highly nonlinear) continuous operators between Banach spaces. The interest of operator learning lies in the fact that it allows to model complex phenomena, e.g. dynamical systems, whose underlying governing equations are not known [DeepOnet, ANIE, NIDE]. The study of operator learning was initiated by the theoretical work [chen], whose implementation was given in [DeepOnet].
The problem of operator learning is therefore to model mappings between infinite dimensional spaces. Since most of the algorithms in practice discretize the domains of the functions, and therefore create a finite space upon which the neural networks are used, one might ask whether the mapping is really between function spaces, or it can just be reduced to some map between high dimensional spaces. For instance, one might ask whether it is possible to upsample the discretized domains after training, e.g. to interpolate and output arbitrarily dense predictions, which would show that the model does not depend on the spacetime stamps used during training. In other words, to what extent the model operates in the infinite dimensional spaces, rather than on some high dimensional Euclidean spaces obtained via discretization.
Various models in operator learning deal with such a question in different ways. For intance, [DeepOnet] fixes points in the domain, but allows continuous inputs. The articles [NODE, ANIE, NIDE] use solvers (for ODEs, IEs or IDEs) to allow continuous sampling/output of points. In [FNO], the use of Fourier transforms allows to upsample the domain of space of functions.
Our perspective in the present work stems from the idea of learning an operator between bases of functions as in [Spectral], where the Chebyshev polynomials have been used for projections as in the Galerkin methods [Fle]. In fact, it was proved in [EZ_proj] that using Leray-Schauder mappings to “nonlinearly project” over a basis of chosen elements approximating a compact set of a Banach space, one can approximate any (possibly nonlinear) operator. This addresses the aforementioned question of learning operators between infinite dimensional spaces rather than just discretization of them, because the approximation happens at the functional level.
In fact, in the present article we introduce a family of neural networks that approximate a compact subseet of a Banach space obtained via Leray-Schauder mappings. Then, the operator learning problem is formulated between the coordinate spaces of the neural networks. This allows us to treat the problem without any discretization, but by using the finite-dimensional reduction provided by the Leray-Schauder mappings to approximate the original learning problem. We show that there exist theoretical guarantees for the approximation to be arbitrarily accurate. In practice, therefore, in this article we also learn the elements upon which the Leray-Schauder mappings project the compact subspace. This represents a paradigmatic shift with respect to the setting of [Spectral], where the projection used is fixed.
We numerically demonstrate the efficacy of the model, and show that it obtains results comparable to state of the art models. The tasks used for comparison are a dataset of integral equations, which shows the efficacy of the model on a task of learning nonlocal operators, and a dataset of Burger’s equation, which explores a PDE problem. The codes for the implementation of this work are found at https://github.com/emazap7/Leray_Schauder_neural_net.
The article is organized as follows. In Section 2 we give some preliminary results and definitions regarding Leray-Schauder mappings and approximations of nonlinear operators. We also prove that it is possible to construct a universal approximator model by using Leray-Schauder mappings and neural networks to model the finite dimensional subspaces obtained via the leray-Schauder mappings. In Section 3 we discuss the details of the deep learning algorithm which implements this theoretical framework. In Section 4, we provide numerical experiments for the algorithms, and compare with other operator learning methods.
2. Theoretical Preliminaries
We recall the following result obtained in [EZ_proj], which gives a method for approximating a continuous (nonlinear) operator between Banach spaces over a compact. The core of the approach is that over a compact it is possible to approximate the domain with finitely many elements, which allows to reformulate the approximation problem in finite dimensions. For this purpose, the Leray-Schauder mappings are used.
Theorem 2.1.
Let $X$ and $Y$ be Banach spaces, let $T:X\longrightarrow Y$ be a continuous (possibly nonlinear) map, and let $K\subset X$ be a compact subset. Then, for any choice of $\epsilon>0$ there exist natural numbers $n,m\in\mathbb{N}$, finite dimensional subspaces $E_{n}\subset X$ and $E_{m}\subset Y$, continuous maps $P_{n}:X\longrightarrow E_{n}$ and $P_{m}:Y\longrightarrow E_{m}$, and a neural network $f_{n,m}:\mathbb{R}^{n}\longrightarrow\mathbb{R}^{m}$ such that for every $x\in K$
(1) $||T(x)-\phi_{m}^{-1}f_{n,m}\phi_{n}P_{n}(x)||<\epsilon,$
where $\phi_{k}:E_{k}\longrightarrow\mathbb{R}^{k}$ indicates an isomorphism between the finite dimensional space $E_{k}$ and $\mathbb{R}^{k}$.
As pointed out in [EZ_proj], the operators $P_{n}$ and $P_{m}$ have an explicit and very simple construction. Following [Topological], let $K$ be a compact, $E_{n}$ an $n$-dimensional space spanned by the elements of an $\epsilon$-net $\{x_{1},\ldots,x_{n}\}$, where $\epsilon>0$ is chosen and fixed. Then $P_{n}:K\longrightarrow E_{n}$ is defined by the assignment
(2) $P_{n}x=\frac{\sum_{i=1}^{n}\mu_{i}(x)x_{i}}{\sum_{j=1}^{n}\mu_{j}(x)},$
where
(3) $\mu_{i}(x)=\begin{cases}\epsilon-\|x-x_{i}\|,\hskip 28.45274pt\|x-x_{i}\|\leq\epsilon\\ 0,\hskip 81.23248pt\|x-x_{i}\|>\epsilon\end{cases}$
for all $i=1,\ldots,n$.
We will refer to these operators as Leray-Schauder projections, although they are not linear, following the same convention as [Topological], which was already adopted in [EZ_proj].
The construction of Theorem 2.1 is based on the fact that an $\epsilon$-net $\{x_{1}.\ldots,x_{n}\}$ is already given. In general, determining such elements of $X$ might be challenging, especially when considering a data space which only approximately represents the actual domain of $T$. When dealing with function spaces, as it is usually the case in real applications, we can approximate the elements $x_{i}$ directly as part of the learning algorithm.
In the present article, our algorithm learns the elements $x_{i}$ used to create the $\epsilon$-net for the Leray-Schauder projection by means of “basis” neural networks. It is therefore natural to ask whether the resulting model, where we learn both the operator in the space $E_{n}$ and the elements whose span give $E_{n}$ is still a universal approximator. The next result, which adapts the arguments of [EZ_proj] to the case of the algorithm in this article, shows that the previous question has a positive answer.
Theorem 2.2.
Let $X$ be the Banach space of continuous functions over a compact $K\subset\mathbb{R}^{N}$ with the uniform norm, $C(K,\mathbb{R}^{M})$. Let $T:X\longrightarrow X$ be a continuous operator, and let $\mathbb{K}\subset X$ be a compact subset. Then, it is possible to find neural networks $\{g_{i}:\mathbb{R}^{N}\longrightarrow\mathbb{R}^{M}\}_{i=1}^{n}$ and $f_{n}:\mathbb{R}^{n}\longrightarrow\mathbb{R}^{n}$ such that for every $x\in\mathbb{K}$
(4) $||T(x)-\phi_{n}^{-1}f_{n}\phi_{n}\tilde{P}_{n}(x)||_{\infty}<\epsilon,$
where $\phi_{k}:F_{k}\longrightarrow\mathbb{R}^{k}$ indicates an isomorphism between the finite dimensional space $F_{k}$ and $\mathbb{R}^{k}$, and $\tilde{P}_{n}$ indicates the Leray-Schauder mapping on the space spanned by the neural networks $g_{i}$.
Proof.
We fix $\epsilon>0$. Using the Heine-Cantor theorem and the construction of Leray-Schauder mappings, it is possible to find functions $x_{i}$, with $i=1,\ldots,n$, such that
(5) $\|x-P_{n}(x)\|_{\infty}$ $<$ $\frac{\epsilon}{6},$
(6) $\|T(z_{1})-T(z_{2})\|_{\infty}$ $<$ $\frac{\epsilon}{3},$
for all $x,z_{1},z_{2}\in\mathbb{K}$. Here the continuous maps $P_{n}$ map $\mathbb{K}$ in the space $E_{n}$ spanned by $x_{i}$, $i=1,\ldots,n$.
The fundamental step of the present proof is to show that we can find neural networks $g_{i}$ spanning a space $F_{n}$ which has the property that $\|x-\tilde{P}_{n}x\|_{\infty}<\frac{\epsilon}{3}$ for all $x\in\mathbb{K}$, where $\tilde{P}_{n}$ is a mapping on $F_{n}$ which we define below. This would allow us to replace the approximation in (5), by an analogous approximation using mappings $\tilde{P}$ and a neural network basis instead of the functions $x_{i}$.
To this purpose, we use the traditional universal approximation results as in [Horn, Fun, LLPS, Lu, Pink] to find neural networks $g_{i}$ such that $\|x_{i}-g_{i}\|_{\infty}$ is arbitrarily small for each $i=1,\ldots,n$, where $x_{i}$ were found above. We define
(7) $\tilde{P}_{n}x=\frac{\sum_{i=1}^{n}\tilde{\mu}_{i}(x)g_{i}}{\sum_{j=1}^{n}\tilde{\mu}_{j}(x)},$
where
$\tilde{\mu}_{i}(x)$ $=$ $\begin{cases}\epsilon-\|x-g_{i}\|_{\infty},\hskip 28.45274pt\|x-g_{i}\|_{\infty}\leq\epsilon\\ 0,\hskip 89.626pt\|x-g_{i}\|_{\infty}>\epsilon.\end{cases}$
In addition, upon taking $g_{i}$ such that $\|x_{i}-g_{i}\|_{\infty}$ is sufficiently small, we can make $|\mu_{i}(x)-\tilde{\mu}_{i}(x)|$ arbitrarily small as well. We prove that
(8) $\inf_{x\in\mathbb{K}}\sum_{j=1}^{n}\mu_{j}(x)>0.$
By way of contradiction, assume otherwise. Then, let $\{y_{k}\}$ be a sequence in $\mathbb{K}$ such that as $k\longrightarrow+\infty$ we have $\sum_{j=1}^{n}\mu_{j}(y_{k})\longrightarrow 0$. Since $\mathbb{K}$ is a compact in a metric space, it is sequentially compact. Let $y_{k_{r}}$ be a convergent subsequence, and let $y_{k_{r}}\longrightarrow y^{*}$. The set $\{x_{1},\ldots,x_{n}\}$ is an $\epsilon$-net by construction, therefore we can find $x_{i}$ such that $q:=\|y^{*}-x_{i}\|_{\infty}<\epsilon$, hence $\epsilon-q>0$. Let $0<\delta<\epsilon-q$. By continuity of the map $\sum_{j=1}^{n}\mu_{i}:\mathbb{K}\longrightarrow\mathbb{R}^{+}_{0}$ we can find a $\delta>0$ such that for all $x\in B(y^{*},\delta)\bigcap\mathbb{K}$ we have $\sum_{j=1}^{n}\mu_{j}(x)\geq\epsilon-q-\delta>0$. Therefore, for $r$ large enough, we have $\sum_{j=1}^{n}\mu_{j}(y_{k_{r}})\geq\epsilon-q-\delta>0$ against the fact that $\sum_{j=1}^{n}\mu_{j}(y_{k})\longrightarrow 0$. This shows that (8) holds true. We put $m:=\inf_{x\in\mathbb{K}}\sum_{j=1}^{n}\mu_{j}(x)$.
Suppose $|\mu_{i}(x)-\tilde{\mu}_{i}(x)|<q$, where $q$ is small enough such that $\sum_{i=1}^{n}\mu_{i}(x)-q>0$. This can be done since $m>0$ as shown above, and it guarantees that $\sum_{j=1}^{n}\tilde{\mu}_{j}(x))$ does not vanish. Then, for all $i=1,\ldots,n$, we have that
$|\frac{\mu_{i}(x)}{\sum_{j=1}^{n}\mu_{j}(x)}-\frac{\tilde{\mu}_{i}(x)}{\sum_{j=1}^{n}\tilde{\mu}_{j}(x)}|$
$\leq$ $\sum_{j=1}^{n}\frac{1}{|(\sum_{j=1}^{n}\mu_{j}(x))(\sum_{j=1}^{n}\tilde{\mu}_{j}(x))|}|\mu_{i}(x)\tilde{\mu}_{j}(x)-\mu_{j}(x)\tilde{\mu}_{j}(x)|$
$\leq$ $\sum_{j=1}^{n}\frac{1}{m(m-q)}[|\mu_{i}(x)\tilde{\mu}_{j}(x)-\mu_{i}(x)\mu_{j}(x)|+|\mu_{i}(x)\mu_{j}(x)-\tilde{\mu}_{i}(x)\mu_{j}(x)|]$
$=$ $\sum_{j=1}^{n}\frac{1}{m(m-q)}[|\mu_{j}(x)|\cdot|\tilde{\mu}_{j}(x)-\mu_{j}(x)|+|\mu_{j}(x)|\cdot|\mu_{i}(x)-\tilde{\mu}_{i}(x)|]$
$<$ $\frac{2n\epsilon q}{m(m-q)},$
which implies that $|\frac{\mu_{i}(x)}{\sum_{j=1}^{n}\mu_{j}(x)}-\frac{\tilde{\mu}_{i}(x)}{\sum_{j=1}^{n}\tilde{\mu}_{j}(x)}|\longrightarrow 0$ as $q\longrightarrow 0$. Since
$|\sum_{i=1}^{n}\frac{\mu_{i}(x)}{\sum_{j=1}^{n}\mu_{j}(x)}-\sum_{i=1}^{n}\frac{\tilde{\mu}_{i}(x)}{\sum_{j=1}^{n}\tilde{\mu}_{j}(x)}|$$\leq$ $\sum_{i=1}^{n}|\frac{\mu_{i}(x)}{\sum_{j=1}^{n}\mu_{j}(x)}-\frac{\tilde{\mu}_{i}(x)}{\sum_{j=1}^{n}\tilde{\mu}_{j}(x)}|$
$<$ $\frac{2n^{2}\epsilon q}{m(m-q)},$
we find that by taking $g_{i}$ such that $\|x_{i}-g_{i}\|_{\infty}$ is small enough, we can obtain $|\sum_{i=1}^{n}\frac{\mu_{i}(x)}{\sum_{j=1}^{n}\mu_{j}(x)}-\sum_{i=1}^{n}\frac{\tilde{\mu}_{i}(x)}{\sum_{j=1}^{n}\tilde{\mu}_{j}(x)}|<\frac{\epsilon}{12\max_{i}\|g_{i}\|_{\infty}}$ and at the same time have $\|x_{i}-g_{i}\|_{\infty}<\frac{\epsilon}{12}$.
Then, in this situation we have
$\|x-\tilde{P}_{n}x\|_{\infty}$
$<$ $\|x-P_{n}x\|_{\infty}+\|P_{n}x-\tilde{P}_{n}x\|_{\infty}$
$=$ $\frac{\epsilon}{6}+\|\frac{\sum_{i}\mu_{i}(x)x_{i}}{\sum_{j}\mu_{j}(x)}-\frac{\sum_{i}\tilde{\mu}_{i}(x)g_{i}}{\sum_{j}\tilde{\mu}_{j}(x)}\|_{\infty}$
$\leq$ $\frac{\epsilon}{6}+\sum_{i=1}^{n}\|\frac{\mu_{i}(x)x_{i}}{\sum_{j=1}^{n}\mu_{j}(x)}-\frac{\tilde{\mu}_{i}(x)g_{i}}{\sum_{j=1}^{n}\tilde{\mu}_{j}(x)}\|_{\infty}$
$\leq$ $\frac{\epsilon}{6}+\sum_{i=1}^{n}[\|\frac{\mu_{i}(x)x_{i}}{\sum_{j=1}^{n}\mu_{j}(x)}-\frac{\mu_{i}(x)g_{i}}{\sum_{j=1}^{n}\mu_{j}(x)}\|_{\infty}+\|\frac{\mu_{i}(x)g_{i}}{\sum_{j=1}^{n}\mu_{j}(x)}-\frac{\tilde{\mu}_{i}(x)g_{i}}{\sum_{j=1}^{n}\tilde{\mu}_{j}(x)}\|_{\infty}]$
$\leq$ $\frac{\epsilon}{6}+\sum_{i=1}^{n}[\frac{\mu_{i}(x)}{\sum_{j=1}^{n}\mu_{j}(x)}\|x_{i}-g_{i}\|_{\infty}+\|g_{i}\|_{\infty}|\frac{\mu_{i}(x)}{\sum_{j=1}^{n}\mu_{j}(x)}-\frac{\tilde{\mu}_{i}(x)}{\sum_{j=1}^{n}\tilde{\mu}_{j}(x)}|]$
$<$ $\frac{\epsilon}{6}+\frac{\epsilon}{12}+\frac{\epsilon}{12}$
$=$ $\frac{\epsilon}{3},$
where we have used the fact that $\mu_{j}(x)\geq 0$ by definition of Leray-Schauder mapping. A similar procedure can be used to find mappings $\tilde{P}_{n^{\prime}}$ for the compact $T(\mathbb{K})$. However, upon possibly passing to subrefinements, we can assume that $n=n^{\prime}$ and $\tilde{P}_{n^{\prime}}=\tilde{P}_{n}$. This applies in the particularly important case when $T(\mathbb{K})\subset\mathbb{K}$.
We let $\phi_{n}$ be an isomorphism which identifies $F_{n}$ with $\mathbb{R}^{n}$, and construct the mapping $T_{n}$ induced by $T$ on $F_{n}$. We find a neural network architecture which approximates $T_{n}$ on the compact $\phi_{n}P_{n}(\mathbb{K})$. With these definitions one can verify that the approximation (4) holds for all $x\in\mathbb{K}$, by adding a bound of type $\|(\phi_{n}^{-1}f_{n}\phi_{n})P_{n}(x)-(\phi_{n}^{-1}f_{n}\phi_{n})\tilde{P}_{n}(x)\|_{\infty}<\epsilon$, using the uniform continuity of $f_{n}$ on the compact $\phi_{n}P_{n}(\mathbb{K})\bigcup\phi_{n}\tilde{P}_{n}(\mathbb{K})$. ∎
The construction of $\tilde{P}_{n}$ is complicated by the fact that we need to map on a space spanned by neural networks through the use of the functions $\tilde{\mu}_{i}$, where we need to find $g_{i}$ such that the maps $\tilde{\mu}_{i}$ have the required properties. This presents a significant implementational, and practical, problem in terms of constructing the mappings $\tilde{P}_{n}$. In practice, we would like to implement $\tilde{P}_{n}$ as neural networks as well, so that the whole architecture consists of neural network components. In the rest of this section we will work toward this direction, showing that it is indeed possible to replace $\tilde{P}_{n}$ by neural networks. This construction will be seen to be a universal approximation architecture on functions that are regular enough in the sense of Hölder. The result is therefore somewhat weaker than the case when one uses $\tilde{P}_{n}$ maps as above. We will denote by $L^{P}_{\mu}([a,b])$ the $L^{p}$-space on $[a,b]$ for some measure $\mu$ (which is assumed to be sufficiently regular). In what follows, we denote the Hölder space by $C^{\kappa,\alpha}([a,b])$, where $\kappa\geq 1$, and $\alpha\in(0,1]$. We assume that $p>2\kappa-1$. We will later mention how to generalize the result to the multivariate case. This relies on cubature formulas, and our choice of proceeding with the single varaible case is motivated by the wish of keeping the notation as simple as possible.
For an interval $[a,b]$ of $\mathbb{R}$, for a discretization (e.g. a partition) $\{t_{1},\ldots,t_{k}\}\subset[a,b]$ and a neural network $F:\mathbb{R}^{k}\longrightarrow\mathbb{R}^{s}$, we define $F(f):=F(f(t_{1}),\ldots,f(t_{k}))$ for any function $f:[a,b]\longrightarrow\mathbb{R}$, and the components of $F$ in $\mathbb{R}^{s}$ are indicated by $F^{i}(f)$. In other words, $F$ is applied to the evaluation of $f$ on the nodes of discretization. In machine learning, one usually has access to data sampled on discretizations of the domains of definition. Therefore, the neural network can work as a proxy of an operator. We leverage this fact to approximate the Leray-Schauder mappings $P_{n}$ via neural networks.
Lemma 2.3.
Let $\mathbb{K}\subset L^{p}_{\mu}([a,b])$ be a compact subspace, and let $P_{n}$ be the Leray-Schauder map $P_{n}:\mathbb{K}\longrightarrow E_{n}$, where $E_{n}$ is spanned by an $\epsilon$-net of smooth functions $\{x_{1},\ldots,x_{n}\}$. Then, for any $\eta>0$, there exists a discretization $\{t_{1},\ldots,t_{k}\}$ of $[a,b]$, and a neural network $F:\mathbb{R}^{\times k}\longrightarrow\mathbb{[}0,+\infty)$ such that
$\|P_{n}(x)-\sum_{i=1}^{n}\frac{F^{i}(x)x_{i}}{\sum_{j=1}^{n}F^{j}(x)}\|_{p}$ $<$ $\eta,$
for all $x\in\mathbb{K}\bigcap\bar{B}_{\rho}$, where $B_{\rho}$ is an open ball in the Hölder sapce $C^{\kappa,\alpha}([a,b])$. for any $\rho>0$ such that $\max_{i}\|x_{i}\|_{\infty}\leq\rho$.
Proof.
We first prove that
(9) $\inf_{x\in\mathbb{K}}\sum_{j=1}^{n}\mu_{j}(x)>0.$
The same argument is found also in the proof of Theorem 2.2, and we report it here in the $L^{p}$ case to render the two proofs readable independently. By way of contradiction, assume otherwise. Then, let $\{y_{k}\}$ be a sequence in $\mathbb{K}$ such that as $k\longrightarrow+\infty$ we have $\sum_{j=1}^{n}\mu_{j}(y_{k})\longrightarrow 0$. Since $\mathbb{K}$ is a compact in a metric space, it is sequentially compact. Let $y_{k_{r}}$ be a convergent subsequence, and let $y_{k_{r}}\longrightarrow y^{*}$. The set $\{x_{1},\ldots,x_{n}\}$ is an $\epsilon$-net by assumption, therefore we can find $x_{i}$ such that $q:=\|y^{*}-x_{i}\|_{p}<\epsilon$, hence $\epsilon-q>0$. Let $0<\delta<\epsilon-q$. By continuity of the map $\sum_{j=1}^{n}\mu_{i}:\mathbb{K}\longrightarrow\mathbb{R}^{+}_{0}$ we can find a $\delta>0$ such that for all $x\in B(y^{*},\delta)\bigcap\mathbb{K}$ we have $\sum_{j=1}^{n}\mu_{j}(x)\geq\epsilon-q-\delta>0$. Therefore, for $r$ large enough, we have $\sum_{j=1}^{n}\mu_{j}(y_{k_{r}})\geq\epsilon-q-\delta>0$ against the fact that $\sum_{j=1}^{n}\mu_{j}(y_{k})\longrightarrow 0$. This shows that (9) holds true. We put $m:=\inf_{x\in\mathbb{K}}\sum_{j=1}^{n}\mu_{j}(x)$. We also set $M:=\sup_{x\in\mathbb{K}}\sum_{j=1}^{n}\mu_{j}(x)$, which is clearly a finite number since each $\mu_{i}$ is continuous over a compact.
[…]
(10) $\hat{\mu}_{i}(u)=\begin{cases}\epsilon-\mathfrak{I}(|x-x_{i}|^{p})^{\frac{1}{p}},\hskip 28.45274pt\mathfrak{I}(|x-x_{i}|^{p})^{\frac{1}{p}}\leq\epsilon\\ 0,\hskip 102.43008pt\mathfrak{I}(|x-x_{i}|^{p})^{\frac{1}{p}}>\epsilon\end{cases}$
for all $i=1,\ldots,n$. We define $\hat{P}_{n}(x):=\sum_{i=1}^{n}\frac{\hat{\mu}_{i}(x)x_{i}}{\sum_{j=1}^{n}\hat{\mu}_{j}(x)}$. Then, for all $x\in\mathbb{K}\bigcap\bar{B}_{\rho}$, whenever $k$ is large enough so that $m-e_{k}>0$, we have
$\|\hat{P}_{n}(x)-P_{n}(x)\|_{p}$$\leq$ $\sum_{i=1}^{n}\|\frac{\hat{\mu}_{i}(x)x_{i}}{\sum_{j=1}^{n}\hat{\mu}_{j}(x)}-\frac{\mu_{i}(x)x_{i}}{\sum_{j=1}^{n}\mu_{j}(x)}\|_{p}$
$=$ $\sum_{i=1}^{n}\|x_{i}\|_{p}\cdot|\frac{\hat{\mu}_{i}(x)}{\sum_{j=1}^{n}\hat{\mu}_{j}(x)}-\frac{\mu_{i}(x)}{\sum_{j=1}^{n}\mu_{j}(x)}|$
$\leq$ $\sum_{i=1}^{n}\frac{\|x_{i}\|_{p}}{(m-e_{k})m}\cdot|\sum_{j=1}^{n}\hat{\mu}_{i}(x)\mu_{j}(x)-\sum_{j=1}^{n}\mu_{i}(x)\hat{\mu}_{j}(x)|$
$\leq$ $\sum_{i=1}^{n}\frac{\|x_{i}\|_{p}}{(m-e_{k})m}\cdot\sum_{j=1}^{n}|\hat{\mu}_{i}(x)\mu_{j}(x)-\mu_{i}(x)\hat{\mu}_{j}(x)|$
$\leq$ $\sum_{i=1}^{n}\frac{\|x_{i}\|_{p}}{(m-e_{k})m}\cdot\sum_{j=1}^{n}[|\hat{\mu}_{i}(x)\mu_{j}(x)-\mu_{i}(x)\mu_{j}(x)|+|\mu_{i}(x)\mu_{j}(x)-\mu_{i}(x)\hat{\mu}_{j}(x)|]$
$=$ $\sum_{i=1}^{n}\frac{\|x_{i}\|_{p}}{(m-e_{k})m}\cdot\sum_{j=1}^{n}[|\mu_{j}(x)|\cdot|\hat{\mu}_{i}(x)-\mu_{i}(x)|+|\mu_{i}(x)|\cdot|\mu_{j}(x)-\hat{\mu}_{j}(x)|]$
$\leq$ $\sum_{i=1}^{n}\frac{\|x_{i}\|_{p}}{(m-e_{k})m}\cdot\sum_{j=1}^{n}Me_{k}$
$=$ $\frac{n^{2}Me_{k}}{m^{2}-me_{k}}\cdot\max_{i}\|x_{i}\|_{p}.$
As we increase the discretization points $k$, i.e. as $k\longrightarrow+\infty$, it follows that $\|\hat{P}_{n}(x)-P_{n}(x)\|_{p}\longrightarrow 0$, and we can therefore approximate $P_{n}$ by means of the map $\hat{P}_{n}$ induced by the quadrature rule $\mathfrak{I}$ over $\mathbb{K}\bigcap\bar{B}_{\rho}$, as long as we choose $k$ large enough. The last step of the proof, is to approximate the quadrature rule by means of a neural network.
Since the functions $\hat{\mu}_{i}:\mathbb{R}^{k}\longrightarrow\mathbb{R}^{+}_{0}$ are continuous with respect to the inputs of the discretization of the integration scheme $\mathfrak{I}$, we can approximate each function $\hat{\mu}_{i}$ by a neural network $F^{i}_{\theta}$, according to the results in [Horn, Fun, LLPS, Lu, Pink] with arbitrarily high precision for each $x\in\mathbb{K}\bigcap\bar{B}_{\rho}$. By thaking the nonlinearity to be a sigmoid, and adding a cutoff at zero we can also ensure that the output of each $F^{i}_{\theta}$ is nonnegative. In particular, we can find architectures such that
$|\frac{\hat{\mu}_{i}(x)}{\sum_{j=1}^{n}\hat{\mu}_{j}(x)}-\frac{F^{i}_{\theta}(x)}{\sum_{j=1}^{n}F^{j}_{\theta}(x)}|<\frac{\eta}{2n\max_{i}\|x_{i}\|_{p}},$
for all $x\in\mathbb{K}\bigcap\bar{B}_{\rho}$. Then, if $k$ is large enough such that $\frac{n^{2}Me_{k}}{m^{2}-me_{k}}\cdot\max_{i}\|x_{i}\|_{p}<\frac{\eta}{2}$, we find that for all $x\in\mathbb{K}\bigcap\bar{B}_{\rho}$ we have
$\|P_{n}(x)-\sum_{i=1}^{n}\frac{F^{i}(x)x_{i}}{\sum_{j=1}^{n}F^{j}(x)}\|_{p}$$\leq$ $\|P_{n}(x)-\hat{P}_{n}(x)\|_{p}+\|\hat{P}_{n}(x)-\sum_{i=1}^{n}\frac{F^{i}(x)x_{i}}{\sum_{j=1}^{n}F^{j}(x)}\|_{p}$
$\leq$ $\frac{n^{2}Me_{k}}{m^{2}-me_{k}}\cdot\max_{i}\|x_{i}\|_{p}+\sum_{i=1}^{n}\|\frac{\hat{\mu}_{i}(x)x_{i}}{\sum_{j=1}^{n}\hat{\mu}_{j}(x)}-\frac{F^{i}_{\theta}(x)x_{i}}{\sum_{j=1}^{n}F^{j}_{\theta}(x)}\|_{p}$
$<$ $\frac{\eta}{2}+\sum_{i=1}^{n}\|x_{i}\|_{p}\cdot|\frac{\hat{\mu}_{i}(x)}{\sum_{j=1}^{n}\hat{\mu}_{j}(x)}-\frac{F^{i}_{\theta}(x)}{\sum_{j=1}^{n}F^{j}_{\theta}(x)}|$
$<$ $\frac{\eta}{2}+\frac{\eta}{2}$
$=$ $\eta.$
This completes the proof, once we stack the neural networks $F^{i}_{\theta}$ to obtain $F=(F^{1}_{\theta},\ldots,F^{n}_{\theta})^{T}$. ∎
Proposition 2.4.
In Lemma 2.3 we can choose the neural network $F$ to consist of a single hidden layer feed forward architecture.
Proof.
For $x\in\mathbb{K}\bigcap\bar{B}_{\rho}$ we have that $\|x\|_{\infty}\leq\rho$. Therefore, for a discretization $\{t_{1},\ldots,t_{k}\}$ we have that $\max_{\ell,i}|x(t_{\ell})-x_{i}(t_{\ell})|\leq 2\rho$. It follows that the inputs of $\mathfrak{I}$ are in the bounded hypercube $[-q,q]^{\times k}$, where $q=2^{p}\rho^{p}$. We can therefore apply the results for universal approximation over compacts. ∎
Theorem 2.5.
Let $T:X\longrightarrow X$ be continuous map, where $X=L^{p}_{\mu}([a,b])$. Let $\mathbb{K}\subset X$ be a compact such that $T(\mathbb{K})\subset\mathbb{K}$. Then, for any given $\epsilon>0$ we can find $n\in\mathbb{N}$, $\rho>0$, and neural networks $f_{n}$ and $F$ such that
(11) $\|T(x)-\phi_{n}^{-1}f_{n}\phi_{n}(\sum_{i=1}^{n}\frac{F^{i}(x)g_{i}}{\sum_{j=1}^{n}F^{j}(x)})\|_{p}$ $<$ $\epsilon,$
for all $x\in\mathbb{K}\bigcap\bar{B}_{\rho}$, where $\bar{B}_{\rho}$ is a radius $\rho$ closed ball in $C^{\kappa,\alpha}([a,b])$, and $g_{i}$ are neural networks for $i=1,\ldots,n$.
Proof.
The proof is similar to that of Theorem 2.2, where the details following Theorem 2.1 do not depend on the use of the uniform norm. The main difference here lies in the use of Lemma 2.3, to construct a neural network Leray-Schauder mapping $F$ which plays the role of $P_{n}$ in Theorem 2.1, and $\tilde{P}_{n}$ in Theorem 2.2. We therefore provide only the details pertaining to this aspect of the proof. Since $T(\mathbb{K})\subset\mathbb{K}$, we can construct a single Leray-Schauder mapping, as opposed to two as in the proof of Theorem 2.1, where one is used for $\mathbb{K}$, and another one for $T(\mathbb{K})$. The double indices referring to the two mappings can therefore be suppressed.
Observe that by density of smooth functions in $X$, we can assume that an $\epsilon$-net is constructed by taking $\{x_{1},\ldots,x_{n}\}$ smooth functions. Furthermore, since we can approximate each $x_{i}$ by neural networks $g_{i}$ in the uniform norm, we can assume that the $\epsilon$-net consists of neural networks. In order to have $g_{i}$ smooth for each $i$, we can restrict ourselves to smooth (nonpolynomial) nonlinearities.
Then, applying Lemma 2.3 we can find $F:\mathbb{R}^{k}\longrightarrow[0,+\infty)$ such that the Leray-Schauder mapping $P_{n}$ can be approximated by the mapping $x\mapsto\sum_{i=1}^{n}\frac{F^{i}(x)g_{i}}{\sum_{j=1}^{n}F^{j}(x)}$ for all $x\in\mathbb{K}\bigcap\bar{B}_{\rho}$ with arbitrary precision, where $\rho\geq\max_{i}\|g_{i}\|_{\infty}$. Now, to apply the arguments of Theorem 2.1 and Theorem 2.2 there is one last subtlety to handle. While $P_{n}(\mathbb{K})$ is a compact in $E_{n}$, it is not a priori clear whether the image of $x\mapsto\sum_{i=1}^{n}\frac{F^{i}(x)g_{i}}{\sum_{j=1}^{n}F^{j}(x)}$, for $x\in\mathbb{K}\bigcap\bar{B}_{\rho}$ is compact. However, by construction we have that
$\|\sum_{i=1}^{n}\frac{F^{i}(x)g_{i}}{\sum_{j=1}^{n}F^{j}(x)}\|_{p}$$\leq$ $\sum_{i=1}^{n}\|\frac{F^{i}(x)g_{i}}{\sum_{j=1}^{n}F^{j}(x)}\|_{p}$
$=$ $\sum_{i=1}^{n}|\frac{F^{i}(x)}{\sum_{j=1}^{n}F^{j}(x)}|\cdot\|g_{i}\|_{p}$
$\leq$ $\sum_{i=1}^{n}|\frac{\epsilon+\eta}{m-\eta}|\cdot\|g_{i}\|_{p}$
$\leq$ $\frac{\epsilon}{m-\eta}\max_{i}\|g_{i}\|_{p},$
where we have denoted by $\eta$ the error in the approximation $P_{n}$ via the mapping $x\mapsto\sum_{i=1}^{n}\frac{F^{i}(x)g_{i}}{\sum_{j=1}^{n}F^{j}(x)}$, used the fact that $\mu_{i}(x)\leq\epsilon$ always, and the properties obtained in the proof of Lemma 2.3. Then, the image of the mapping $x\mapsto\sum_{i=1}^{n}\frac{F^{i}(x)g_{i}}{\sum_{j=1}^{n}F^{j}(x)}$ is bounded in the finite dimensional space $E_{n}$, and we can therefore bound it by a closed ball in $E_{n}$ which is compact, since $E_{n}$ with the norm induced by the $p$ norm of $X$ is equivalent to the $n$-dimensional Euclidean space. So, the neural network $f_{n}$ can be defined over a compact, where it is uniformly continuous. Therefore, we can choose the $\eta$ approximation of $P_{n}$ by $x\mapsto\sum_{i=1}^{n}\frac{F^{i}(x)g_{i}}{\sum_{j=1}^{n}F^{j}(x)}$ such that $\|f_{n}(\phi_{n}P_{n}(x))-f_{n}\phi_{n}\sum_{i=1}^{n}\frac{F^{i}(x)g_{i}}{\sum_{j=1}^{n}F^{j}(x)}\|_{p}<\frac{\epsilon}{\|\phi_{n}^{-1}\|}$. Then, we have
$\|T(x)-(\phi_{n}^{-1}f_{n}\phi_{n})\sum_{i=1}^{n}\frac{F^{i}(x)g_{i}}{\sum_{j=1}^{n}F^{j}(x)}\|_{p}$$\leq$ $\|T(x)-TP_{n}(x)\|_{p}+\|TP_{n}(x)-P_{n}TP_{n}(x)\|_{p}$
$+\|P_{n}TP_{n}(x)-(\phi_{n}^{-1}f_{n}\phi_{n})P_{n}(x)\|_{p}$
$+\|(\phi_{n}^{-1}f_{n}\phi_{n})P_{n}(x)-(\phi_{n}^{-1}f_{n}\phi_{n})\sum_{i=1}^{n}\frac{F^{i}(x)g_{i}}{\sum_{j=1}^{n}F^{j}(x)}\|_{p}$
$<$ $\frac{\epsilon}{3}+\frac{\epsilon}{3}+\frac{\epsilon}{3}+\epsilon$
$=$ $2\epsilon.$
Since $\epsilon$ was arbitrary, the proof is complete. ∎
Remark 2.6.
To consider the multivariate case, one needs to use the Hölder space of functions with multiple variables, and needs to substitute cubatures to quadrature rules. Sobolev cubature formulas [Sobolev] can then be used to control the errors in the approximations of the integrations, and as the number of points grow, one can approximate the multidimensional integrals with arbitrary precision.
3. Algorithm
Theorem 2.1 gives a foundation for the algorithm discussed in this section, whose implementation is the main objective of the present article. Our deep learning algorithm is based on constructing a function input from the data available for initialization (e.g. an initial condition or boundary values, depending on the formulation of the problem), and (nonlinearly) project the input on a basis of functions modeled with neural networks. The projection maps are Leray-Schauder maps, defined through Equation (1). The maps $\mu_{i}$ are either fixed as in Equation (3) (for some choice of norm), or are learned as in Theorem 2.2 and Theorem 2.5. In fact, while our experiments show that the assignment in Equation (3) gives a functioning model, we have found that hyperparameter fine tuning is more complex in this setting, and simply learning the $\mu_{i}$ maps, which in turn means that we learn the Leray-Schauder maps as well, makes training much simpler.
The model consists of $n$ feed forward neural networks $g_{i}$, $i=1,\ldots,n$, where each $g_{i}:K\longrightarrow\mathbb{R}^{M}$, and a feed forward neural network $f:\mathbb{R}^{n}\longrightarrow\mathbb{R}^{n}$. Since we also learn the functions $\mu_{i}$ in Equation (3), we also have convolutional neural networks $\mu_{i}:C(K,\mathbb{R}^{M})\longrightarrow\mathbb{R}^{+}_{0}$. In practice, we use CNNs to implement the $\mu_{i}$ neural networks because they mimic an integration procedure which can be performed in higher dimensions. They take a discretized input function and produce a numerical value (which can be constrained to be nonnegative by choice of the final nonlinearity).
The neural networks $g_{i}$ have the role of projection basis, since they are the functions whose span gives the finite dimensional approximation space in the proof of Theorem 2.5.
We use the initialization values (data available during inference) to create a function $\mathbf{y}(\mathbf{x},t)$, where $\mathbf{x}\in C\subset\mathbb{R}^{n}$, and $t\in[a,b]$, where $K=C\times[a,b]$ in the setting of Theorem 2.2. This is the input to the operator. In practice, we obtain $\mathbf{y}$ by interpolating the input data, e.g. $\mathbf{y}(\mathbf{x},0)$ and $\mathbf{y}(\mathbf{x},1)$ in the experiments below. Then, we compute the Leray-Schauder mapping as
$P_{n}(\mathbf{y})=\sum_{i=1}^{n}\frac{\mu^{i}(\mathbf{y})g_{i}}{\sum_{j=1}^{n}\mu^{j}(\mathbf{y})},$
which is the same operator used in the approximation of Theorem 2.5, where $F^{i}$ has been named $\mu^{i}$ to mimic the definition in (3). Since $P_{n}(\mathbf{y})$ can be evaluated through the networks $g_{i}$, it is clear that $P_{n}$ takes a function as input, obtained via initialization data, and produces a function as output. This function $P_{n}(\mathbf{y})$ lives inside the span $\langle g_{1},\ldots,g_{n}\rangle$. In fact, $P_{n}(\mathbf{y})$ is a linear combination of $g_{i}$, with coefficients $q_{i}:=\frac{\mu^{i}(\mathbf{y})}{\sum_{j=1}^{n}\mu^{j}(\mathbf{y})}$. The vector of coefficients $\mathbf{q}=(q_{1},\ldots,q_{n})^{T}\in\mathbb{R}^{n}$ is therefore the input of the network $f$, following the proof of Theorem 2.5.
We compute $\mathbf{b}=f(\mathbf{q})$, and $\mathbf{b}$ is used to take a linear combination of the neural networks $g_{i}$, which we set $\psi=\sum_{i=1}^{n}b_{i}g_{i}$. We can then evaluate $\psi(\mathbf{x},t)$ and compute the loss $\|\psi(\mathbf{x},t)-\mathbf{z}(\mathbf{x},t)\|$, where $\mathbf{z}$ is the target data which we are predicting.
This is summarized in Algorithm 1. We will refer to this method as the Leray-Schauder Neural Operator for brevity.
Algorithm 1 Algorithm for the Leray-Schauder neural operator.
1:Neural networks $\{g_{i}\}_{i=1}^{n}$ and $f_{n}$
2:Approximation $\phi_{n}^{-1}f_{n}\phi_{n}\tilde{P}_{n}$ of operator $T$
3:Use initialization data to define input function $y$ (e.g. linear interpolation)
4:Project $y$ on $\langle g_{i}\rangle$ using $\tilde{P}$: Coefficients $q_{i}$
5:Use $\phi_{n}$ to identify coefficients $\mathbf{q}$ with elements of $\mathbb{R}^{n}$
6:Apply $f_{n}$ on $\mathbf{q}$: $f(\mathbf{q})=\mathbf{b}$
7:Output function, using $\phi^{-1}_{n}$, is $\psi(\mathbf{x},t)=\sum_{i=1}^{n}b_{i}g_{i}(\mathbf{x},t)$
8:Compute loss with target data $\mathbf{z}$: $\|\psi-\mathbf{z}\|$
9:Compute gradients and use stochastic grandient descent to optimize
4. Experiments
To demonstrate the capabilities of the method discussed in this article we experiment on two datasets. One consists of a set of spirals generated by solving an integral equation, and the other is a dataset on Burgers’ equation. In both cases, the model has access at initialization to initial time $t=0$ and final time $t=1$ (initial and final time). The initialization is obtained by linearly interpolating between initial and final time configurations.
Table 1.
IE Spirals Burgers’
Original Interpolation $s=256$ $s=512$
Leray-Schauder $0.0011\pm 0.0005$ $0.0011\pm 0.0005$ $0.0017\pm 0.0012$ $0.0017\pm 0.0011$
ANIE $0.0014\pm 0.0002$ $0.0015\pm 0.0003$ $0.0016\pm 0.0016$ $0.0017\pm 0.0018$
Spectral NIE $0.0022\pm 0.0016$ $0.0028\pm 0.0032$
FNO1D (init 5) $0.0292\pm 0.0285$ $0.0994\pm 0.1207$
FNO1D (init 10) $0.0286\pm 0.0268$ $0.1541\pm 0.2724$
Both datasets were used in [ANIE, Spectral], and are publicly available at https://figshare.com/articles/dataset/IE_spirals/25606242 and https://figshare.com/articles/dataset/Burgers_1k_t400/25606149.
The $2D$ integral equation spirals have been obtained by numerically solving an IE via iterative methods (Banach-Caccioppoli iterations). The solver is our implementation found in [ANIE], and whose code is publicly available. The equation has the form:
[…]
where $z_{0}$ was sampled from a uniform distribution to derive different instances of the equation for the dataset.
For the Burgers’ equation, our dataset is generated using the Matlab code used in [FNO], which can be found in their GitHub page: https://github.com/zongyi-li/fourier\_neural\_operator/tree/master/data\_generation/burgers. The solution is given on a spatial mesh of $1024$ and $400$ time points are generated from a random initial condition. We use $1000$ curves for training and test on $200$ unseen curves. Initial and final times are seen by the model at initialization, and all the time points constitute the ground truth for training the model. Some details regarding the Burgers’ equation are given in [ANIE].
The implementational details of our model are found on our GitHub page https://github.com/emazap7/Leray_Schauder_neural_net.
In the experiments, we find that the model is comparable with the state of the art model (ANIE, [ANIE]) on both datasets. For the IE spirals, we have inlcluded two experiments, one on prediction of the original dynamics, and another one on an interpolation task where the model has access to half the points during training and predicts all of them during evaluation. For the Burgers’ dataset we have experimented on two spatial resolutions $s=256$ and $s=512$. An example of ground truth dynamics for the Burgers’ equation is found in Figure 1, while the corresponding model’s prediction is in Figure 2. The vertical direction indicates space and the horizontal direction indicates time.
Figure 1. Example of ground truth data for Burgers’ dynamics where $\vec{x}$-axis is time and $\vec{y}$-axis represents space
Figure 2. Example of model’s prediction for Burgers’ dynamics where $\vec{x}$-axis is time and $\vec{y}$-axis represents space
We notice that in the spirals tasks, FNO1D is initialized using more points ($5$ or $10$ intead of $2$ as in the Leray-Schauder neural operator). Therefore, FNO1D is given an advantage in initialization. For the Burgers’ equation, FNO2D was seen to perform significantly worse than ANIE in [ANIE], and we have therefore compared only with the latter in the present article for a more concise treatment.